# Memory-Enhanced Neural Solvers for Efficient Adaptation in Combinatorial Optimization

### v2

@@ Abstract
Combinatorial Optimization is crucial to numerous real-world applications, yet still presents challenges due to its (NP-)hard nature. Amongst existing approaches, heuristics often offer the best trade-off between quality and scalability, making them suitable for industrial use. While Reinforcement Learning (RL) offers a flexible framework for designing heuristics, its adoption over handcrafted heuristics remains incomplete within industrial solvers. Existing learned methods still lack the ability to adapt to specific instances and fully leverage the available computational budget. The current best methods either rely on a collection of pre-trained policies, or on data-inefficient fine-tuning; hence failing to fully utilize newly available information within the constraints of the budget. In response, we present MEMENTO, an approach that leverages memory to improve the adaptation of neural solvers at inference time. MEMENTO enables updating the action distribution dynamically based on the outcome of previous decisions. We validate its effectiveness on benchmark problems, in particular Traveling Salesman and Capacitated Vehicle Routing, demonstrating its superiority over tree-search and policy-gradient fine-tuning; and showing it can be zero-shot combined with diversity-based solvers. We successfully train all RL auto-regressive solvers on large instances, and show that MEMENTO can scale and is data-efficient. Overall, MEMENTO enables to push the state-of-the-art on $11$ out of $12$ evaluated tasks.
****footnotetext: Work completed during an internship at InstaDeep
@@ 1 Introduction
Combinatorial Optimization (CO) encompasses a vast range of real-world applications, ranging from transportation (J-F Audy & Rousseau, 2011) and logistics (Dincbas et al., 1992) to energy management (Froger et al., 2016). These problems involve finding optimal orderings, labels, or subsets of discrete sets to optimize given objective functions. Real-world CO problems are typically NP-hard with a solution space growing exponentially with the problem size, making it intractable to find the optimal solution. Hence, industrial solvers rely on sophisticated heuristic approaches to solve them in practice. Reinforcement Learning (RL) provides a versatile framework for learning such heuristics and has demonstrated remarkable success in tackling CO tasks (Mazyavkina et al., 2021).
@@
Traditionally, RL methods train policies to incrementally construct solutions. However, achieving optimality in a single construction attempt for NP-hard problems is impractical. Therefore, pre-trained policies are often combined with search procedures. The literature introduces a range of these procedures from stochastic sampling (Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023), beam search (Steinbiss et al., 1994; Choo et al., 2022), Monte Carlo Tree Search (MCTS) (Browne et al., 2012) to online fine-tuning (Bello et al., 2016; Hottung et al., 2022) and searching a space of diverse pre-trained policies (Chalumeau et al., 2023b). One popular online fine-tuning strategy, Efficient Active Search (EAS) (Hottung et al., 2022), uses transitions from generated solutions to derive policy gradient updates. However, it suffers from the inherent drawbacks of back-propagation, in particular having each data point only impacting the update as much as what is enabled by the learning rate. The leading strategy, COMPASS (Chalumeau et al., 2023b), relies on a space of diverse pre-trained policies for fast adaptation, but is limited to selecting the most appropriate one, and lacks an update mechanism using the collected data. Meanwhile, the use of memory has grown in modern deep learning, demonstrated by the success of retrieval augmented approaches in Natural Language Processing (Lewis et al., 2020), and to a smaller extent, its use in RL (Humphreys et al., 2022). These mechanisms create a closer link between collected experience and policy update, making them a promising candidate to improve adaptation.
@@
In this vein, we introduce MEMENTO, a method to efficiently update the action distribution of neural solvers online, leveraging a memory of recently collected data. Importantly, MEMENTO is model agnostic and can be applied to many existing neural solvers. MEMENTO is able to learn expressive update rules, which prove experimentally to outperform classic policy gradient updates. We evaluate our method on two popular routing problems - Traveling Salesman (TSP) and Capacitated Vehicle Routing (CVRP). We evaluate all methods both in and out of distribution, and tackle instances up to size 500 to better understand their scaling properties. We experimentally show that MEMENTO can be combined with existing methods and provide systematic improvements, reaching state-of-the-art on $11$ out of $12$ tasks. We provide an analysis of the update rules learned by MEMENTO, examining how it enables to adapt faster than policy gradient methods like EAS.
@@
Our contributions come as follows: (i) We introduce MEMENTO, a flexible and scalable method – composed of a memory and a processing module – enabling efficient adaptation of policies at inference time. (ii) We provide experimental evidence that MEMENTO can be combined with existing approaches to boost their performance in and out of distribution, even for large instances and unseen solvers. (iii) Whilst doing so, we train and evaluate leading construction methods on TSP and CVRP instances of size $500$ solely with RL, proving to outperform all existing RL methods, and release the checkpoints. (iv) We open-source the implementation of MEMENTO in JAX (Bradbury et al., 2018), along with test sets and checkpoints to facilitate future research advancements.
@@ 2 Related work
Construction approaches in RL for CO refer to methods that incrementally build a solution one action after the other. Hopfield & Tank (1985) was the first to use a neural network to solve the TSP, followed by Bello et al. (2016) and Deudon et al. (2018) who respectively added a reinforcement learning loss and an attention-based encoder. These works were further extended by Kool et al. (2019) and Kwon et al. (2020) to use a general transformer architecture (Vaswani et al., 2017), which has become the standard model choice that we also leverage in this paper. These works have given rise to several variants, improving the architecture (Xin et al., 2021; Luo et al., 2023) or the loss (Kim et al., 2021; 2022; Drakulic et al., 2023; Sun et al., 2024). These improvements are orthogonal to our work and could a priori be combined with our method. It is to be noted that construction approaches are not restricted to routing problems: numerous works have tackled various CO problems, especially on graphs, like Maximum Cut (Dai et al., 2017; Barrett et al., 2020), or Job Shop Scheduling Problem (JSSP) (Zhang et al., 2020; Park et al., 2021).
@@
The first type of approach focuses on improving the diversity of the generated solutions (Kwon et al., 2020; Grinsztajn et al., 2023; Chalumeau et al., 2023b; Hottung et al., 2024). POMO (Kwon et al., 2020) makes use of different starting points to enable the same policy to generate diverse candidates. Poppy (Grinsztajn et al., 2023) leverages a population of agents with a loss targeted at specialization on sub-distribution of instances. It was extended in Chalumeau et al. (2023b) and Hottung et al. (2024), respectively replacing the population with a continuous latent space (COMPASS) or a discrete context vector (PolyNet).
@@
The second category of methods, which includes MEMENTO, addresses the improvement aspect. These methods are theoretically orthogonal and can be combined with those mentioned earlier. EAS (Hottung et al., 2022) employs a parameter-efficient fine-tuning approach on the test instances, whereas SGBS (Choo et al., 2022) enhances this strategy by incorporating tree search. While these two methods demonstrate impressive performance, they rely on rigid, handcrafted improvement and search procedures, which may not be optimal for diverse problem domains or computational budgets.
In contrast, MOCO, FER and MOCO (Dernedde et al., 2024; Jingwen et al., 2023; I. Garmendia et al., 2024) learn the policy improvement update. MOCO (Dernedde et al., 2024) introduces a meta-optimizer that learns to calculate flexible parameter updates based on the reinforce gradient, remaining budget, and best solutions discovered so far. MARCO (I. Garmendia et al., 2024) aggregates graph edges’ features and introduce problem-dependent similarity metrics to improve exploration of the solution space. Both approaches enable to learn adaptation strategies but are tied to heatmap-based policies, preventing their applicability to certain CO problems like CVRP. FER (Jingwen et al., 2023) learns a rule to update the instance nodes’ embedding, whereas MEMENTO directly updates the action logits, hence being architecture-agnostic, scaling better and achieving higher performance.
@@ 3.1 Preliminaries
The specificity of most practical cases in RL for CO is that the policy is given a budget of shots $B$ to find the best possible solution, rather than a single shot. Consequently, the learning objective should rather be : $\pi^{*}=\operatorname*{arg\,max}\nolimits_{\pi}\mathbb{E}_{\tau_{i}\sim\pi}[\max\nolimits_{i=1,...,B}R(\tau_{i})]$. It is hence beneficial to learn policies with higher entropy to provide better exploration, and ideally, policies that can use their previous attempts to (i) explore decisions they have not tried yet (ii) put more focus on promising regions of the space.
@@ 3.2 MEMENTO
Figure 1: MEMENTO uses a memory to adapt neural solvers at inference time. When taking a decision, data from similar states is retrieved and prepared (1,2), then processed by a MLP to derive correction logits for each action (3). Summing the original and new logits enables to update the action distribution. The resulting policy is then rolled out (4), and transitions’ data is stored in a memory (5,6), including node visited, action taken, log probability, and obtained return.
A compelling approach is to store all past attempts in a memory which can be leveraged for subsequent trajectories. This ensures that no information is lost and that promising trajectories can be used more than once. There are many ways of implementing such framework depending on how the information is stored, retrieved, and used in the policy. We would like the memory-based update mechanism to be (i) learnable (how to use past trajectories to craft better trajectories should be learnt instead of harcoded) (ii) light-weight to not compromise inference time unduly (iii) agnostic to the underlying model architecture (iv) able to leverage pre-trained memory-less policies.
@@
To this end, we introduce MEMENTO, a method that dynamically adapts the action distribution of the neural solver using online collected data. This approach, illustrated on Fig. 1, achieves the four desiderata: it stores light-weight data about past attempts, and uses an auxiliary model to update the action logits directly based on the previous outcomes. This allows MEMENTO to be used on top of any pretrained policy and to be totally agnostic to its architecture. In principle, MEMENTO can be combined with most existing construction RL algorithms: Attention-Model (Kool et al., 2019), POMO (Kwon et al., 2020), Poppy  (Grinsztajn et al., 2023), COMPASS (Chalumeau et al., 2023b), which we confirm experimentally in Section 4.2. It learns an update rule during training, enabling data-efficient adaptation compared to policy gradient methods. The data retrieved from the memory and used by the auxiliary model to compute the new action logits contains more than the usual information used to derive a policy gradient update, for instance the budget remaining, enabling to calibrate the exploration/exploitation trade-off and discover superior update rules. We show in Appendix F that MEMENTO has capacity to at least rediscover REINFORCE, we provide a comparative plot on Fig. 4 and we show empirically in Section 4.1 that we outperform the leading policy-gradient updates adaptation method.
@@
In the memory, we store data about our past attempts. Akin to a replay buffer, this data needs to reflect past decisions, their outcome, and must enable to take better future decisions. In our memory, for each transition experienced while constructing a solution, we store the following: (i) node visited (ii) action taken, corresponding to the node that the policy decided to visit next (iii) log-probability given to that action by the model (iv) return of the entire trajectory (negative cost of the solution built) (v) budget at the time that solution was built (vi) log-probability that was suggested by the memory for that action (vii) log-probability associated with the entire trajectory and (viii) the log-probability associated with the remaining part of the trajectory. We can observe that elements (ii, iii and iv) are those needed to reproduce a REINFORCE update, and other features are additional context that will help for credit assignment and distribution shift estimation. Note that the remaining budget will also be added to the data when deriving the policy update.
@@
Each time an action must be taken in a given state, we want to retrieve data in the memory that can inform us on past decisions that were taken in a similar situation. Since this retrieval process will happen at each step, we do not want it to be too costly. A good speed/relevance balance is achieved by retrieving data collected in the same node we are currently in; being faster than k-nearest neighbor retrieval, while extracting data of similar relevance.
@@
Once data has been retrieved, it is used to update the action logits of the base policy. From the data, we separate the actions from their associated features (logp, return, etc…). We concatenate the remaining budget to the features. Each feature is normalised, and the resulting feature vector is processed by a Multilayer Perceptron (MLP) $H_{\theta_{M}}$ which outputs a scalar weight. Each action is one-hot encoded, and a weighted average of the actions vectors is computed based on the scalar weight obtained. This aggregation outputs a new vector of action logits. This vector of logits is summed with the vector of logits output by the base model. The following paragraph introduces the mathematical formalism, and Fig. 2 illustrates it.
More formally, when visiting a node, with a given state $s$, we retrieve from the memory data experienced in the past when visiting that same node. This data, $M_{s}$, is a sequence of tuples $(a_{i},f_{a_{i}})$, where $a_{i}$ are the past actions tried and $f_{a_{i}}$ are various features associated with the corresponding trajectories, as discussed above. The update is computed by $l_{M}=\sum_{i}\mathbf{a_{i}}H_{\theta_{M}}(f_{a_{i}})$, where $\mathbf{a_{i}}$ is the one-hot encoding of action $a_{i}$. Let $l$ be the logits from the base policy, the final logits used to sample the next action are given by $l+l_{M}$. We show in Appendix F that in the worst case, this enables MEMENTO to re-discover the REINFORCE update.
@@
Existing construction methods are trained for one-shot optimization, with a few exceptions that are trained for few-shots (Grinsztajn et al., 2023; Chalumeau et al., 2023b; Hottung et al., 2024). We adapt the training process to take into account the budgeted multi-shot setting in which methods are actually used in practice. We hence rollout the policy as many times as the budget allows on the same instance before taking an update. Our loss is inspired by ECO-DQN (Barrett et al., 2020) and ECORD (Barrett et al., 2022): for a new given trajectory, the reward is the ReLU of the difference between the return and the best return achieved so far on that instance. Hence, those summed rewards equal the best score found over the budget. In practice, to account for the fact that it is harder to get an improvement as we get closer to the end of the budget, we multiply this loss with a weight that increases logarithmically as the budget is consumed. See Appendix B for explicit formula and Appendix B for pseudo-code.
@@
At the end of the training process, we obtain the parameters of the memory processing module, which can be used with any neural solver with no back-propagation needed anymore. The parameters of the base neural solver and the weights of the memory processing module are fixed, and the adaptation rules while simply be derived by filling the memory with the collected attempts and by processing the data retrieved from the memory. All hyperparameters can be found in Appendix C.

### v3

# Memory-Enhanced Neural Solvers for Routing Problems

@@ Abstract
Routing Problems are central to many real-world applications, yet remain challenging due to their (NP-)hard nature. Amongst existing approaches, heuristics often offer the best trade-off between quality and scalability, making them suitable for industrial use. While Reinforcement Learning (RL) offers a flexible framework for designing heuristics, its adoption over handcrafted heuristics remains incomplete. Existing learned methods still lack the ability to adapt to specific instances and fully leverage the available computational budget. Current best methods either rely on a collection of pre-trained policies, or on RL fine-tuning; hence failing to fully utilize newly available information within the constraints of the budget. In response, we present MEMENTO, an approach that leverages memory to improve the search of neural solvers at inference. MEMENTO leverages online data collected across repeated attempts to dynamically adjust the action distribution based on the outcome of previous decisions. We validate its effectiveness on the Traveling Salesman and Capacitated Vehicle Routing problems, demonstrating its superiority over tree-search and policy-gradient fine-tuning; and showing that it can be zero-shot combined with diversity-based solvers. We successfully train all RL auto-regressive solvers on large instances, and verify MEMENTO’s scalability and data-efficiency: pushing the state-of-the-art on $11$ out of $12$ evaluated tasks.
Figure 1: MEMENTO outperforms Efficient Active Search on a wide range of instance sizes and different evaluation budgets. Its time complexity makes it scalable to large instances and policies.
[…]
@@ 1 Introduction
Combinatorial Optimization (CO) encompasses a vast range of applications, ranging from transportation (J-F Audy and Rousseau, 2011) and logistics (Dincbas et al., 1992) to energy management (Froger et al., 2016). These problems involve finding optimal orderings or labels to optimize given objective functions. Real-world CO problems are typically NP-hard with a solution space growing exponentially with the problem size, making it intractable to find the optimal solution. Hence, industrial solvers rely on sophisticated heuristic approaches to solve them in practice. Reinforcement Learning (RL) provides a versatile framework for learning such heuristics and has demonstrated remarkable success in tackling CO tasks (Mazyavkina et al., 2021).
@@
Traditionally, RL methods train policies to incrementally construct solutions. However, achieving optimality in a single attempt for NP-hard problems is impractical. Therefore, pre-trained policies are often combined with search procedures. The literature introduces a range of these procedures, from stochastic sampling (Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023), beam search (Steinbiss et al., 1994; Choo et al., 2022), Monte Carlo Tree Search (MCTS) (Browne et al., 2012) to online fine-tuning (Bello et al., 2016; Hottung et al., 2022) and searching a space of diverse pre-trained policies (Chalumeau et al., 2023b). One popular online fine-tuning strategy, Efficient Active Search (eas) (Hottung et al., 2022), uses transitions from generated solutions to derive policy gradient updates. However, it suffers from the inherent drawbacks of back-propagation, in particular having each data point only impacting the update as much as what is enabled by the learning rate. The leading strategy, compass (Chalumeau et al., 2023b), relies on a space of diverse pre-trained policies for fast adaptation, but is limited to selecting the most appropriate one, and lacks an update mechanism using the collected data. Meanwhile, the use of memory has grown in modern deep learning, demonstrated by the success of retrieval augmented approaches in Natural Language Processing (Lewis et al., 2020), and to a smaller extent, its use in RL (Humphreys et al., 2022). These mechanisms create a closer link between collected experience and policy update, making them a promising candidate to improve adaptation.
@@
In this vein, we introduce memento, a method to update the action distribution of neural solvers online, leveraging a memory of recently collected data. The approach is model agnostic and can be applied to existing neural solvers. memento learns expressive update rules, which experimentally prove to outperform standard policy-gradient updates. We evaluate our method on two popular routing tasks, the Traveling Salesman Problem (TSP) and the Capacitated Vehicle Routing Problem (CVRP). We evaluate all methods both in and out-of-distribution and tackle instances up to size 500 to better understand their scaling properties. We provide an analysis of the update rules learned by memento, examining how it enables to adapt faster than policy gradient methods like eas.
@@
Our contributions come as follows: (i) We introduce memento, a memory and a processing module, enabling efficient adaptation of policies at inference time. (ii) We provide experimental evidence that memento can be combined with existing approaches to boost performance in and out-of-distribution, even for large instances and unseen solvers, reaching state-of-the-art (SOTA) on $11$ out of $12$ tasks. (iii) Whilst doing so, we train and evaluate leading construction methods on TSP and CVRP instances of size $500$ solely with RL, outperforming all existing RL methods. (iv) We open-source our implementations in JAX (Bradbury et al., 2018), along with test sets and checkpoints.
@@ 2 Related work
These refer to methods that incrementally build a solution one action after the other. Hopfield and Tank (1985) was the first to use a neural network to solve TSP, followed by Bello et al. (2016) and Deudon et al. (2018) who respectively added an RL loss and an attention-based encoder. These works were further extended by Kool et al. (2019) and Kwon et al. (2020) to use a general transformer architecture (Vaswani et al., 2017), which has become the standard model choice that we also leverage in this paper. These works have given rise to several variants, improving the architecture (Xin et al., 2021; Luo et al., 2023) or the loss (Kim et al., 2021, 2022; Drakulic et al., 2023; Sun et al., 2024). These improvements are orthogonal to our work and could a priori be combined with our method. Construction approaches are not restricted to routing problems: numerous works have tackled various CO problems, especially on graphs, like Maximum Cut (Dai et al., 2017; Barrett et al., 2020), or Job Shop Scheduling Problem (JSSP) (Zhang et al., 2020; Park et al., 2021).
@@
The first type of approach focuses on improving the diversity of the generated solutions (Kwon et al., 2020; Grinsztajn et al., 2023; Chalumeau et al., 2023b; Hottung et al., 2024). pomo (Kwon et al., 2020) uses different starting points to enable the same policy to generate diverse candidates. poppy (Grinsztajn et al., 2023) leverages a population of agents with a loss targeted at specialization on sub-distribution of instances. It was extended in Chalumeau et al. (2023b) and Hottung et al. (2024), respectively replacing the population with a continuous latent space (compass) or a discrete context vector (PolyNet).
@@
The second category of methods, which includes memento, addresses the improvement aspect. These methods are theoretically orthogonal and can be combined with those mentioned earlier. eas (Hottung et al., 2022) and Meta-SAGE (Son et al., 2023) employ a parameter-efficient fine-tuning approach on the test instances, whereas sgbs (Choo et al., 2022) enhances this strategy by incorporating tree search. While demonstrating good performance, they rely on rigid, handcrafted improvement and search procedures, which may not be optimal for diverse problems and computational budgets.
In contrast, several methods aim to learn these improvement mechanisms. Macfarlane et al. (2022) train a Graph Neural Network to perform a tree search akin to MCTS. moco, fer, and marco (Dernedde et al., 2024; Jingwen et al., 2023; I. Garmendia et al., 2024) learn the policy improvement update. moco (Dernedde et al., 2024) introduces a meta-optimizer that learns to calculate flexible parameter updates based on the reinforce gradient, remaining budget, and best solutions discovered so far. marco (I. Garmendia et al., 2024) aggregates graph edges’ features and introduce problem-dependent similarity metrics to improve exploration of the solution space. Both approaches enable to learn adaptation strategies but are tied to heatmap-based policies, preventing their applicability to certain CO problems like CVRP. fer (Jingwen et al., 2023) learns a rule to update the instance nodes’ embedding, whereas memento directly updates the action logits, hence being architecture-agnostic.
@@ 3.1 Preliminaries
The specificity of most practical cases in RL for CO is that the policy is given a number of attempts allowed per instance (budget $B$), to find the best possible solution, rather than a single attempt. Consequently, the learning objective should rather be : $\pi^{*}=\operatorname*{arg\,max}\nolimits_{\pi}\mathbb{E}_{\tau\sim\pi}[\max\nolimits_{i=1,...,B}R(\tau_{i})]$.
@@ 3.2 MEMENTO
Figure 2: MEMENTO uses a memory to adapt neural solvers at inference time. When taking a decision, data from similar states is retrieved and prepared (1,2), then processed by a MLP to derive correction logits for each action (3). Summing the original and new logits enables to update the action distribution. The resulting policy is then rolled out (4), and transitions’ data is stored in a memory (5,6), including node visited, action taken, log probability, and return obtained.
A compelling approach is to store all past attempts in a memory that can be leveraged for subsequent trajectories. This ensures that no information is lost and that promising trajectories can be used more than once. There are many ways of implementing such framework depending on how the information is stored, retrieved, and used in the policy. We would like the memory-based update mechanism to be (i) learnable (how to use past trajectories to craft better trajectories should be learnt instead of harcoded) (ii) light-weight to not compromise inference time unduly (iii) agnostic to the underlying model architecture (iv) able to leverage pre-trained memory-less policies.
@@
To this end, we introduce memento, a method that dynamically adapts the action distribution of the neural solver using online collected data. This approach, illustrated on Fig.˜2, achieves the four desiderata: it stores light-weight data about past attempts, and uses an auxiliary model to update the action logits based on the previous outcomes. It can be used on top of any pre-trained policy (architecture-agnostic). In principle, memento can be combined with existing construction RL algorithms: Attention-Model (Kool et al., 2019), pomo (Kwon et al., 2020), poppy (Grinsztajn et al., 2023), compass (Chalumeau et al., 2023b), which we confirm experimentally. It learns an update rule during training, enabling data-efficient adaptation compared to policy gradient methods. The data retrieved from the memory and used by the auxiliary model to compute the new action logits contains more than the typical information used to derive a policy gradient update. For instance, the budget remaining, enabling to calibrate the exploration/exploitation trade-off and discover superior update rules. We show in Appendix˜F that memento has capacity to at least rediscover REINFORCE, and we show empirically in Section˜4.1 that we outperform the leading policy-gradient updates adaptation method.
@@
In the memory, we store data about past attempts (Fig.˜2, step 6). Akin to a replay buffer, this data needs to reflect past decisions, their outcome, and must enable to take better future decisions. For each transition experienced while constructing a solution, the memory stores the following: (i) node visited (i.e., in the context of TSP or CVRP, this corresponds to the current city or customer location), (ii) action taken, corresponding to the node that the policy decided to visit next (iii) log-probability given to that action by the model (iv) return of the entire trajectory (negative cost of the solution built) (v) budget at the time that solution was built (vi) log-probability that was suggested by the memory for that action (vii) log-probability associated with the entire trajectory and (viii) the log-probability associated with the remaining part of the trajectory. We can observe that elements (ii, iii and iv) are those needed to reproduce a REINFORCE update, and other features are additional context that will help for credit assignment and distribution shift estimation. The remaining budget is also added to the data when deriving the policy update.
@@
Each time an action is taken in a given state, we retrieve data from the memory as shown in step 1 of Fig.˜2. We retrieve data that inform the policy on past decisions that were taken in similar situations. To achieve a good balance between speed and relevance, we retrieve data collected in the same node as we are currently in; showing to be significantly faster than k-nearest neighbour retrieval, while extracting data of similar relevance. More details and motivation for this design choice provided in Appendix˜I.
@@
Once data has been retrieved, it is used to derive a policy update. We separate the actions from their associated features (logp, return, etc…). We concatenate the remaining budget to the features. Each feature is normalised, and the resulting feature vector (Fig.˜2, step 2) is processed by a Multilayer Perceptron (MLP) $H_{\theta_{M}}$ which outputs a scalar weight. We compute the logits from the MLP output (Fig.˜2, step 3): each action is one-hot encoded, and a weighted average of the actions vectors is computed based on the scalar weight obtained. This aggregation outputs a new vector of action logits. We sum the new vector of logits with the vector of logits output by the base model. The following paragraph introduces the mathematical formalism, and Fig.˜3 illustrates it schematically.
More formally, when visiting a node, with a given state $s$, we retrieve from the memory data experienced in the past when visiting that same node. This data, $M_{s}$, is a sequence of tuples $(a_{i},f_{a_{i}})$, where $a_{i}$ are the past actions tried and $f_{a_{i}}$ are various features associated with the corresponding trajectories, as discussed above. The update is computed by $l_{M}=\sum_{i}\mathbf{a_{i}}H_{\theta_{M}}(f_{a_{i}})$, where $\mathbf{a_{i}}$ is the one-hot encoding of action $a_{i}$. Let $l$ be the logits from the base policy, the final logits used to sample the next action are given by $l+l_{M}$. We show in Appendix˜F that memento has enough capacity to re-discover the REINFORCE update.
@@
Existing construction methods are trained for one-shot optimization, with a few exceptions trained for few-shots (Grinsztajn et al., 2023; Chalumeau et al., 2023b; Hottung et al., 2024). We adapt the training process to fit the multi-shot setting in which methods are used in practice. We rollout the policy (Fig.˜2, step 4-5) as many times as the budget allows on the same instance before taking an update. Our loss is inspired by ECO-DQN (Barrett et al., 2020) and ECORD (Barrett et al., 2022): for a new given trajectory, the reward is the ReLU of the difference between the return and the best return achieved so far. Hence, those summed rewards equal the best score found over the budget. In practice, to account for the fact that it is harder to get an improvement as we get closer to the end of the budget, we multiply this loss with a weight that increases logarithmically as the budget is consumed. See Appendix˜B for explicit formula and pseudo-code.
@@
The memory processing module can be used with any neural solver. The parameters of the base neural solver and those of the memory processing module are frozen (no more backpropagation), and the action updates are derived by filling the memory with the collected attempts and by processing the data retrieved from the memory, as described in the previous paragraphs, until the inference budget is consumed. All hyperparameters can be found in Appendix˜C.